# Predicted third-order sweet spots for $\phi_{0}$-junction Josephson parametric amplifiers

### v1

@@ I Introduction
Quantum limited parametric amplifiers are key components for low-noise microwave signal amplification in qubit readout circuits [1, 2, 3]. The most common parametric amplifiers are based on the so-called fourth order nonlinear term of the inductive potential, also known as the Kerr term. This term provides for robust and practical operation. However, it leads to pump-dependent frequency shifts, constrains the pump frequency to values close to the signal frequency, and also causes gain saturation at higher signal powers, limiting the dynamic range of amplification [4, 5].
@@
We propose that a skewed $\phi_{0}$-junction can be used to realize a SNAIL-like device based on a single Josephson junction [42]. The inductive potential of a nanowire $\phi_{0}$-junction can be optimized to zero-out fourth-order Kerr terms. We find the Kerr-free sweet spots for two-harmonic and three-harmonic CPRs derived from tight-binding simulations, and compared with experiments.
By way of assessment of device characteristics from key metrics such as power gain, 1 dB compression power and operable frequency bandwidth, Josephson junctions based on Sn as superconductor offer the critical current levels (100s of nA) that are compatible with SNAIL operation. The device requires large magnetic fields in order to break time-reversal symmetry. Such fields are present by default in certain quantum applications such as quantum dot-based spin qubits and putative toplogical qubits. For near-zero field use, fixed local fields can be generated by micromagnets. Single nanowire parametric amplifiers can potentially be more compact and simpler to control. Additionally, gate-voltage tunability of nanowire junctions facilitates control over a large bandwidth of working frequency (on the order of a few GHz).
@@ II Skewed $\phi_{0}$-junction
Figure 1: (a) Amplitudes $I_{n}$ from numerical simulations as a function of field $B$. (b) Numerical data (symbols) and fits (solid lines) for $I_{n}=\eta_{n1}(1-\eta_{n2}B^{2})$, where $\eta_{n1}$ and $\eta_{n2}\approx 1$ are fitting parameters. (c) Skewed $\phi_{0}$-junction potential as function of field $B$ and phase $\phi$ for optimized values $a=8.83$ and $c=9.60$. (d) Line-cuts of potential $U(\phi)$ from panel (c).
@@
This function lacks symmetry upon inversion of both $\phi$ and $I$, allowing for large third-order nonlinearities in the Josephson potential. In quasi one-dimensional weak links, we assume that amplitudes $I_{1}$ and $I_{2}$, as well as phase shifts $\phi_{0}$ and $\delta_{12}$ are dependent on magnetic field $B$ applied along the spin orbit field direction:
$\begin{split}I(\phi)&=\alpha(1-B^{2})\sin(\phi+aB)\\ &+\beta(1-B^{2})\sin(2\phi+2aB+cB)\end{split}$ (2)
where $\phi_{0}=aB$, $\delta_{12}=cB$, $I_{1}=\alpha(1-B^{2})$, $I_{2}=\beta(1-B^{2})$, for $B<<B_{c}$, where $B_{c}=1$ is the critical field. $a$, $c$, $\alpha$ and $\beta$ are constants. The Josephson potential is:
$\begin{split}U(\phi)=&-\alpha(1-B^{2})\cos(\phi+aB)\\ &-\frac{\beta}{2}(1-B^{2})\cos(2\phi+2aB+cB)\end{split}$ (3)
@@
For the amplitudes of the Josephson harmonics $I_{n}$ we use values from the Fourier expansion of the numerical current-phase relation [46]. In Fig. 1(a) we plot $I_{n}$ vs $B$ from simulation data. In Fig. 1(b), we show the fit for the first three harmonics of the CPR. The fit functions closely follow the postulated quadratic relation. We also find that phases $\phi_{0}$ and $\delta_{12}$ increase linearly with field within the range $\sim 0-100$ mT, here we set $B_{c}=200$mT for fitting purposes. We find the ratio $I_{1}/I_{2}\sim 4$. We perform our analysis in the next section using the first two Josephson harmonics, and we expand to a third in the Appendix B.
@@
Figure 2: Non-linear coefficients $c_{n}$ as function of $B$ for a set of optimized values $a=8.83$ and $c=9.60$. Sweet spots with$|c_{4}|\propto 0$ and a significant $|c_{3}|$ are at $B\approx\pm 0.26,\pm 0.39$ shown by red and black dashed lines.
@@
The sweet spots are not exclusive to these specific values of $a$ and $c$. In fact, there can be multiple such points on a pareto frontier. This increases the likelihood of reaching these settings in real nanowire Josephson junctions where CPR can vary from device to device. This is illustrated in Fig. 3, where we fix $a$ and vary $c$ while monitoring $c_{3}$ and $c_{4}$. The condition for $c_{4}=0$ and $c_{3}\neq 0$ is met over a large parameter space. Also, varying $a$ for a fixed $c$ does not change $c_{3}$ or $c_{4}$ because $a$ parametrizes the global phase shift $\phi_{0}$, this is shown in the Appendix A.
In the above calculations, we consider only the first two harmonics of the CPR. However, the dynamics in a real device may be influenced by the presence of higher-order harmonic terms. In Appendix B, we show how adding a third order harmonic in the CPR would affect the nonlinear behavior of the device. While the behavior is more complex, the Kerr-free sweet spots are still present.
@@ IV Device proposal
Figure 4: (a) Schematic of a parametric amplifier circuit with nanowire (NW) Josephson junction coupled to a resonator. A micromagnet (MM) is placed close to the junction and NW electron density is tuned with a gate voltage $V_{g}$. (b) Detailed schematic of the NW and MM with thin superconducting shell (S), a break in the shell defines the JJ. An external magnetic field $B_{\text{ext}}$ can be used in conjunction with the micromagnet to fine-tune the effective field near the sweet spot.
The $\phi_{0}$-junction Josephson parametric amplifier device consists of a hybrid superconductor-semiconductor nanowire junction as the key nonlinear component incorporated into a resonator and coupled to a transmission line (Fig. 4(a)). As discussed above, the nonlinearity can be tuned with a magnetic field applied in the effective spin-orbit field direction, typically perpendicular to the nanowire, to induce the skewed and asymmetric CPR effect. If zero-field operation is of interest, we propose using stray fields from a micromagnet placed in proximity to the nanowire (Fig. 4(b)). In previous studies [48, 49, 50, 51] micromagnets made of materials such as CoFe were used to generate strong local fields (upto±$\sim\pm 100$ mT) near the nanowire junction. Junctions with thin superconducting shells have high in-plane critical fields of the order of $\sim 1$ T [52]. A small parallel field can also be applied to fine-tune the total magnetic field to a Kerr-free sweet spot. The field strength and direction can be designed using micromagnetics simulations platforms such as MuMax3 [53]. By placing the large superconducting shapes a distance away from the micromagnet, the global circuit including the resonator can be unaffected by local magnetic field effects: we estimate microTelsa stray perpendicular fields at distances of tens of microns away from the micromagnet.
@@
In order to estimate the behavior of the $\phi_{0}$-junction in a microwave circuit, we calculate the experimentally relevant terms $g_{3}$ and $g_{4}$ for the SNAIL Hamiltonian that includes the JJ and the resonator. Under rotating wave approximation and $\phi\ll 1$ (detailed in Appendix C), the Hamiltonian is:
@@
where $\omega_{0}$ is the bare resonator frequency, $L_{J}$ is the Josephson junction inductance and $L$ is the linear inductance of the circuit. The characteristic frequency $\omega_{r}=\frac{\omega_{0}}{\sqrt{(}1+L_{J}/c_{2}L)}$  [10] should fall within a few GHz frequency range typical for standard microwave experiments as shown in Fig. 11. For Sn-InSb and Sn-InAs hybrid nanowire junctions, critical current at zero field is of order $100-500$ nA, tunable with gate voltage. Therefore, the Josephson energy $E_{J}\sim 50-250$ GHz. Choosing realistic circuit parameters we apply the magnetic field evolution discussed in Sec. II to estimate $g_{3}$ and $g_{4}$ (Fig. 5).
@@
Following Fig. 5, at the Kerr-free point $|g_{3}|/2\pi\approx 33$ MHz and $\omega_{r}/2\pi\approx 3.98$ GHz. The term $g_{4}$ is negligible ($|g_{4}|/2\pi\approx 12$ Hz limited by the numerical resolution). Assuming $|\alpha_{s}|\approx|\alpha_{i}|=0.01$ for low signal power and taking the coupling rate $\kappa/2\pi=0.4$ GHz, we find a gain of $G_{0}=20$ dB that can be achieved over a dynamic bandwidth of $40$ MHz, shown in the Appendix D.
The input signal power for which the gain drops by 1dB, $P_{-1\text{dB}}$, is typically dependent on $|g_{4}|$ for fourth-order amplifiers. In SNAILs, $g_{3}$ as well as other factors can lead to gain saturation [10]. For optimal performance $g_{3}$ may need to be reduced in addition to tuning $\kappa$ and the energy participation ratio [54].
@@ V Conclusions
We find sweet spots where third-order nonlinearity dominates while the fourth-order nonlinearity is suppressed, an ideal condition for implementing the three-wave mixing scheme. We show that these sweet spots are present across a wide range of parameters of the two-component CPR. The sweet spots persist when a third-order Josephson harmonic is added to the CPR. We estimate the device performance to meet the standards of state-of-the-art quantum-limited parametric amplifiers. We also predict that the electrostatic gate tuning of the NW JJ will enable a wide range of operational frequencies for amplification and potential three-wave mixing. We propose a micromagnet-based design for near-zero field operation. The single-junction approach can help miniaturize three-wave Josephson parametric amplifiers which sometimes employ $\sim$20 SNAIL arrays [10]. Future work will focus on the experimental implementation of this device, including the design and the optimization of the microwave circuit implemented using magnetic-field compatible superconducting materials.
@@ VIII Acknowledgments
Appendix A Dependence on a global phase shift $\phi_{0}$
@@
Appendix B Sweet spots with higher order CPR harmonics
@@
We also look at how the addition of the third harmonic phase shift $\delta_{13}$ influences the characteristic behavior of $c_{n}$. We show $c_{n}$ as a function of $d$ and $B$ at $a=1$, $c=5$ in Fig. 8. As the relative strength of the third harmonic component $I_{3}$ is small, the potential landscape is not changed significantly. However, in the special case of $I_{2}$ and $I_{3}$ being dominantly large, it may perturb the behavior as discussed in the Appendix B.1.
Finally, we also look at the higher order nonlinear terms and search for points where all nonlinear terms $c_{n}\approx 0$ except for $c_{5}$ and $c_{6}$, shown in Fig.  9. In this case, at $a=1$, $c=20$ we find points where $c_{6}$ is the dominant term at $B\sim\pm 0.6$, $d\sim\pm 5$. More such points exist at higher values of $c$ and across a range of values of $d$. It has been suggested in the literature that the higher order nonlinear interactions can be useful for hardware-efficient quantum error correction schemes [55, 42].
@@
B.1 CPR with large second and third harmonic components
@@
Appendix C Nonlinear Hamiltonian
@@
Upon Hamiltion quantization, where the phase operator is $\hat{\phi}=\phi_{\text{zpf}}\left(\hat{a}^{\dagger}+\hat{a}\right)$ and the number operator is $\hat{N}=iN_{\text{zpf}}\left(\hat{a}^{\dagger}-\hat{a}\right)$ (zpf stands for zero point fluctuation), $\hat{a}^{\dagger}$ and $\hat{a}$ are the bosonic creation and annihilation operators of the harmonic oscillator, such that $\left[\hat{\phi},\hat{N}\right]=i$, $[\hat{a},\hat{a}^{\dagger}]=1$, and $\phi_{\text{zpf}}=\left(\frac{2E_{C}}{c_{2}E_{J}}\right)^{1/4}$ where $c_{n}=\frac{1}{n!}\frac{\partial^{n}U}{\partial\phi}\big{|}_{\phi=\tilde{\phi}_{\text{min}}}$[56, 57, 54], the total Hamiltonian takes the form:
@@
Appendix D Amplifier performance estimation
Figure 12: (a) Power gain vs. pump amplitude. (b) Power gain vs. signal frequency at optimum pump amplitude.
At the sweet spot, $|g_{3}|\approx 33$ MHz, $g_{4}\approx 12$ Hz and $\omega_{r}\approx 3.98$ GHz. Using these parameters, we estimate the power gain of $\phi_{0}$-junction Josephson parametric amplifier for the three-wave mixing scheme. Using Eq. (6) and assuming $|\alpha_{s}|\approx|\alpha_{i}|=0.01$ for low signal power and $\kappa=400$ MHz, we find an optimum pump strength for achieving maximum gain shown in Fig.  12(a). In Fig. 12(b), we show gain vs. signal strength $|\alpha_{s}|$ with maximum gain at $\omega_{s}=\omega_{p}/2$, and a gain of $20$ dB achieved over a bandwidth of $40$ MHz. These values are an upper estimate on performance, which in real devices may be limited by additional factors that are not included in this calculation.

### v2

@@ I Introduction
Quantum limited parametric amplifiers are key components for low-noise microwave signal amplification in qubit readout circuits [44, 53, 3]. The most common parametric amplifiers are based on the so-called fourth order nonlinear term of the anharmonic potential, also known as the Kerr term. This term provides for robust and practical operation. However, it leads to pump-dependent frequency shifts, constrains the pump frequency to values close to the signal frequency, and also causes gain saturation at higher signal powers, limiting the dynamic range of amplification [16, 25].
@@
We propose that a skewed $\phi_{0}$-junction can be used to realize a SNAIL-like device based on a single Josephson junction [46]. The anharmonic potential of a nanowire $\phi_{0}$-junction can be optimized to zero-out fourth-order Kerr terms. We find the Kerr-free sweet spots for two-harmonic and three-harmonic CPRs derived from tight-binding simulations, and compared with experiments.
Josephson junctions based on Sn as superconductor offer the critical current levels (100s of nA) that are compatible with SNAIL operation. The device requires large magnetic fields in order to break time-reversal symmetry. Such fields are present by default in certain quantum applications such as quantum dot-based spin qubits and putative toplogical qubits. For near-zero field use, fixed local fields can be generated by micromagnets. Single nanowire parametric amplifiers can potentially be more compact and simpler to control. Additionally, gate-voltage tunability of nanowire junctions facilitates control over a large bandwidth of working frequency (on the order of a few GHz).
@@ II Skewed $\phi_{0}$-junction
Figure 1: (a) Amplitudes $I_{n}$ from numerical simulations as a function of field $B$. (b) Numerical data (symbols) and fits (solid lines) for $I_{n}=\alpha_{n}(1-\beta_{n}B^{2})$, where $\alpha_{n}$ and $\beta_{n}$ are fitting parameters. (c) Skewed $\phi_{0}$-junction potential as function of field $B$ and phase $\phi$ for values $a=8.83$ and $c=9.60$ at an optimal point. (d) Line-cuts of potential $U(\phi)$ from panel (c).
@@
where $I_{1}$, $I_{2}$ are the amplitudes of the first and second harmonic components, $\phi_{0}$ is a global phase shift and $\delta_{12}$ is the relative phase shift between the first and second harmonic components. Relative phase $\delta_{12}$ contributes to the skew in the current phase relation which can be quantified as:
$\eta=\frac{|I_{c+}|-|I_{c-}|}{|I_{c+}|-|I_{c-}|}$ (2)
where $I_{c+}$ and $I_{c-}$ are the critical currents in the positive and negative directions, and when $I_{2}\ll I_{1}$, $\eta=-2I_{2}/I_{1}\sin(\delta_{12})$ [38].
The current phase relation for a skewed $\phi_{0}$ junction can be derived from a simple microscopic model of a single channel nanowire with Rashba spin orbit coupling, in the presence of an external magnetic field. The so called ’diode effect’ or the skewed $\phi_{0}$ behavior comes from interplay of spin-orbit coupling along with the external magnetic field causing two interacting psuedospin bands in the weak link Andreev bound states. Using this model [38], it has been shown that a CPR of this form can be obtained in both short and long junction limits. The skew can be tuned with external magnetic field following the analysis in this paper, where strength of $\eta$ varies under the short vs long junction limit depending on the effect of external field in the leads and in the junction. The general CPR in this case has multiple higher harmonic components, however when the first two harmonics are dominant the CPR is approximated as Eq. (1). The tunability of $\delta_{12}$ with external magnetic field has been discussed extensively in [54] both using simulations and experimental data.
The function in Eq. (1) lacks symmetry upon inversion of both $\phi$ and $I$, allowing for large third-order nonlinearities in the Josephson potential. In quasi one-dimensional weak links, we assume that amplitudes $I_{1}$ and $I_{2}$, as well as phase shifts $\phi_{0}$ and $\delta_{12}$ are dependent on magnetic field $B$ applied along the spin orbit field direction:
$\begin{split}I(\phi,B)&=\alpha_{1}(1-B^{2})\sin(\phi+aB)\\ &+\alpha_{2}(1-B^{2})\sin(2\phi+2aB+cB)\end{split}$ (3)
where $\phi_{0}=aB$, $\delta_{12}=cB$, $I_{1}=\alpha_{1}(1-B^{2})$, $I_{2}=\alpha_{2}(1-B^{2})$, for $B<<B_{c}$, where $B_{c}=1$ is the critical field. $a$, $c$, $\alpha_{n}$ are constants.
The Josephson potential is:
$\begin{split}U(\phi,B)=&-\alpha_{1}(1-B^{2})\cos(\phi+aB)\\ &-\frac{\alpha_{2}}{2}(1-B^{2})\cos(2\phi+2aB+cB)\end{split}$ (4)
@@
For the amplitudes of the Josephson harmonics $I_{n}$ we use values from the Fourier expansion of the numerical current-phase relation [54]. In Fig. 1(a) we plot $I_{n}$ vs $B$ from simulation data. In Fig. 1(b), we show the fit for the first three harmonics of the CPR. The fit functions closely follow the postulated quadratic relation. We find $\alpha_{1}/\alpha_{2}\sim 4$, $\beta_{n}\sim 1$ when $B\ll B_{c}$ ($\beta_{1}\sim 0.9$ and $\beta_{2}\sim 1.4$ for $B_{c}=100$ mT). Beyond the limit $B\ll B_{c}$, the effective $B_{c}$ varies for each higher harmonic $I_{n}$. For further calculations in this paper, we consider a fixed $B_{c}$ for all $I_{n}$, i.e., take $\beta=1$ following the minimal model for CPR as shown in Eq. (1).
Phases $\phi_{0}$ and $\delta_{12}$ depend linearly on field when $B\ll B_{c}$, as shown in Fig. (6) in Appendix A. Using the ratio $I_{1}/I_{2}\sim 4$, we perform our analysis in the next section using the first two Josephson harmonics, and we expand to a third in the Appendix C.
@@
[…]
@@
The sweet spots are not exclusive to these specific values of $a$ and $c$. In fact, there can be multiple such points on a pareto frontier. This increases the likelihood of reaching these settings in real nanowire Josephson junctions where CPR can vary from device to device. This is illustrated in Fig. 3, where we fix $a$ and vary $c$ while monitoring $c_{3}$ and $c_{4}$. The condition for $c_{4}=0$ and $c_{3}\neq 0$ is met over a large parameter space. Also, varying $a$ for a fixed $c$ does not change $c_{3}$ or $c_{4}$ because $a$ parametrizes the global phase shift $\phi_{0}$, this is shown in the Appendix B.
In the above calculations, we consider only the first two harmonics of the CPR. However, the dynamics in a real device may be influenced by the presence of higher-order harmonic terms. In Appendix C, we show how adding a third order harmonic in the CPR would affect the nonlinear behavior of the device. While the behavior is more complex, the Kerr-free sweet spots are still present.
@@ IV Device proposal
Figure 4: (a) Schematic of a parametric amplifier circuit with nanowire (NW) Josephson junction coupled to a resonator. A micromagnet ($M$) is placed close to the junction and NW electron density is tuned with a gate voltage $V_{g}$. (b) Detailed schematic of the NW and $M$ with thin superconducting shell (S), a break in the shell defines the $J$. An external magnetic field $B_{\text{ext}}$ can be used in conjunction with the micromagnet to fine-tune the effective field near the sweet spot.
[…]
@@
In order to estimate the behavior of the $\phi_{0}$-junction in a microwave circuit, we calculate the experimentally relevant terms $g_{3}$ and $g_{4}$ for the SNAIL Hamiltonian that includes the JJ and the resonator. Under rotating wave approximation and $\phi\ll 1$ (detailed in Appendix D), the Hamiltonian is:
@@
where $\omega_{0}$ is the bare resonator frequency, $L_{J}$ is the Josephson junction inductance and $L$ is the linear inductance of the circuit. The characteristic frequency $\omega_{r}=\frac{\omega_{0}}{\sqrt{(}1+L_{J}/c_{2}L)}$  [19] should fall within a few GHz frequency range typical for standard microwave experiments as shown in Fig. 12. For Sn-InSb and Sn-InAs hybrid nanowire junctions, critical current at zero field is of order 100500100-500 nA, tunable with gate voltage. Therefore, the Josephson energy $E_{J}\sim 50-250$ GHz. Choosing realistic circuit parameters we apply the magnetic field evolution discussed in Sec. II to estimate $g_{3}$ and $g_{4}$ (Fig. 5).
@@
Following Fig. 5, at the Kerr-free point $|g_{3}|/2\pi\approx 32$ MHz and $\omega_{r}/2\pi\approx 3.9$ GHz. The term $g_{4}$ is negligible ($|g_{4}|/2\pi\approx 2-0.012$ kHz limited by the numerical resolution). Assuming $|\alpha_{s}|\approx|\alpha_{i}|=0.01$ for low signal power and taking the coupling rate $\kappa/2\pi=0.4$ GHz, we find a gain of $G_{0}=20$ dB can be achieved over a bandwidth of $40$ MHz, shown in the Appendix E.
The input signal power for which the gain drops by 1 dB, $P_{-1\text{dB}}$, is typically dependent on $|g_{4}|$ for fourth-order amplifiers. In SNAILs, $|g_{3}|$ in addition to other factors can also lead to gain saturation due to pump depletion  [19]. For optimal performance $g_{3}$ may need to be reduced in addition to tuning $\kappa$ and the energy participation ratio [39]. The estimated values for $P_{\text{-1dB}}$ of $\phi_{0}$ junction parametric amplifier is expected to be on par with the state of the art parametric amplifiers, detailed discussion provided in Appendix E.
@@ V Conclusions
We find sweet spots where third-order nonlinearity dominates while the fourth-order nonlinearity is suppressed, an ideal condition for implementing the three-wave mixing scheme. We show that these sweet spots are present across a wide range of parameters of the two-component CPR. The sweet spots persist when a third-order Josephson harmonic is added to the CPR. We estimate the device performance to meet the standards of state-of-the-art quantum-limited parametric amplifiers. We also predict that the electrostatic gate tuning of the NW JJ will enable a wide range of operational frequencies for amplification and potential three-wave mixing.
In [54], using experimental data and KWANT simulation analysis, it has been shown that skew patterns appear at multiple gate voltages. Thus varying the weak link chemical potential using a gate voltage could be used to tune the position of the sweet point vs magnetic field and the effective strength of $|g_{3}|/|g_{4}|$. However, as seen from the data, there is no clear trend for skewness ($\eta$) as a function of gate voltage. And since there is no analytical or phenomenological model for the behavior, it hasn’t been included in the scope of this study.
We propose a micromagnet-based design for near-zero field operation. The single-junction approach can help miniaturize three-wave Josephson parametric amplifiers which sometimes employ $\sim$20 SNAIL arrays [19]. Future work will focus on the experimental implementation of this device, including the design and the optimization of the microwave circuit implemented using magnetic-field compatible superconducting materials.
@@ VIII Acknowledgments
Appendix A Relative phase $\delta_{12}$ vs. $B$
The relative phase $\delta_{12}$ has a linear dependence on external magnetic field $B$ when $B\ll B_{c}$ as shown in Fig. (6), where $\phi_{i0}$ is the global phase of $i$th harmonic in CPR. This data has been taken from KWANT simulation results in [54].
Figure 6: (a) Global phase shift $\phi_{i0}$ of $i$th harmonic in CPR as a function of external field $B$ [54]. (b) $\delta_{12}$ is linearly dependent on $B$ for $B\ll B_{c}$.
Appendix B Dependence on a global phase shift $\phi_{0}$
@@
Appendix C Sweet spots with higher order CPR harmonics
@@
We also look at how the addition of the third harmonic phase shift $\delta_{13}$ influences the characteristic behavior of $c_{n}$. We show $c_{n}$ as a function of $d$ and $B$ at $a=1$, $c=5$ in Fig. 9. As the relative strength of the third harmonic component $I_{3}$ is small, the potential landscape is not changed significantly. However, in the special case of $I_{2}$ and $I_{3}$ being dominantly large, it may perturb the behavior as discussed in the Appendix C.1.
Finally, we also look at the higher order nonlinear terms and search for points where all nonlinear terms $c_{n}\approx 0$ except for $c_{5}$ and $c_{6}$, shown in Fig.  10. In this case, at $a=1$, $c=20$ we find points where $c_{6}$ is the dominant term at $B\sim\pm 0.6$, $d\sim\pm 5$. More such points exist at higher values of $c$ and across a range of values of $d$. It has been suggested in the literature that the higher order nonlinear interactions can be useful for hardware-efficient quantum error correction schemes [40, 46].
@@
C.1 CPR with large second and third harmonic components
@@
Appendix D Nonlinear Hamiltonian
@@
[…]
@@
Appendix E Amplifier performance estimation
Figure 13: (a) Power gain vs. pump amplitude. (b) Power gain vs. signal frequency and pump amplitude. (c) Power gain vs. signal frequency and signal amplitude at $|\alpha_{p}|=1.5$
At the sweet spot $|g_{3}|/2\pi\approx 32$ MHz, $\omega_{r}\approx 3.9$ GHz, and $|g_{4}|/2\pi$ can range from a few kHz to few Hz ($\approx 2$ kHz - $12$ Hz) depending on numerical resolution of optimization. Here we present a theoretical estimation based on ideal performance of the device.
[…]
Following [19], we can estimate $P_{\text{-1dB}}$ as:
$P_{\text{-1dB}}^{\text{Stark}}~\frac{\kappa}{|g_{4}|}\frac{1}{G_{0}^{5/4}}\hbar\omega_{r}\kappa\sim-84\text{dB}$ (15)
At the sweet point, assuming $|g_{4}|\approx 0$, the gain is limited by pump depletion given as:
$P_{\text{-1dB}}^{\text{pump dep}}~\frac{\kappa}{g_{3}^{2}/\omega_{r}}\frac{1}{G_{0}^{3/2}}\hbar\omega_{r}\kappa\sim-110\text{dB}$ (16)
where $|g_{3}|/2\pi=32.2$ MHz, $|g_{4}|/2\pi=1.94$ kHz, $\omega_{r}/2\pi=3.9$ GHz at the sweet point, $|\alpha_{s}|\approx|\alpha_{i}|=0.01$, $\kappa/2\pi=400$ MHz, $G_{0}=20$ dB.
In this case, the gain is limited by saturation due to pump depletion when operating at the sweet point. Here $g_{3}$, $g_{4}$ values are subjected to numerical limitations. However, there would be other practical limitations contributing to the maximum gain achieved, which is beyond the scope of our numerical analysis. For example, nanowire-to-nanowire variability in non-linearity and critical current, variability based on device design and materials, and effect of magnetic field on the readout circuit and coupling to the nanowire junction, etc.